(1 point) Perform the following calculation over Z3 R:#f8-03-16
Answer #1

Given But 41 (mod 3) 6 = 0 (mod 3) in 3. → [ 1-63] = P -?] But -2=1[mod 3) -1 = 2 (mod 3) in 723 - [1 2]